# Dissociation of Faithful and Unfaithful Reasoning in LLMs

### v1

@@ Appendix C Sensitivity Analysis
C.1 Experiment 1
@@
Figure 19 shows the complete hallucination rates against the faithful recovery rates for experiment 1. Consistent with our analysis from Section 5, we find that complete hallucination shows distinct behavior compared tofaithful recovery (and the overall error recovery trend), from copy errors to calculation errors ($p<0.005$) or from calculation errors to propagated calculation errors ($p<0.001$).
C.2 Experiment 2
Figure 20: Sensitivity analysis for faithful vs. unfaithful recoveries. Difference in GPT-4’s recovery rates between between large error and small error conditions. Negative values indicate recoveries occurred more often for small errors. Error bars are 95% confidence intervals.
In experiment 2, we examined the effect of error magnitude on GPT-4’s faithful and unfaithful error recovery rates. From Figure 20, we see that GPT-4 generally performed complete hallucination less frequently for large errors than small errors, whereas it performed faithful recovery more frequently for large errors than small errors. As in Section 6, there is a significant difference between the responses of faithful and unfaithful recovery rates to error size ($p<0.001$).
C.3 Experiment 3
In experiment 3, we measured the effect of context noise on error recovery rates. As in the primary analysis, context noise increases overall recovery rates as well as faithful recoveries. There is also a significant difference in the effect on faithful vs. unfaithful recoveries ($p<0.001$). However, the sensitivity analysis did not identify an effect on unfaithful recovery rates.
Figure 21: Sensitivity analysis for faithful vs. unfaithful recoveries. Difference in GPT-4’s recovery rates between context noise and baseline conditions. Negative values indicate that recoveries occurred more often in the baseline condition. Error bars indicate 95% confidence intervals.

### v2

@@ Appendix D Sensitivity Analysis
D.1 Experiment 1
Figure 17: Sensitivity analysis for faithful vs. unfaithful recoveries, examining the difference in recovery rates between between large error and small error conditions. Negative values indicate recoveries occurred more often for small errors. Error bars are 95% confidence intervals.
In experiment 1, we examined the effect of error magnitude on the faithful and unfaithful error recovery rates of GPT-4, Claude-3 Opus, and Llama-3 70B. From Figure 17, we see that GPT-4 generally performed complete hallucination less frequently for large errors than small errors, whereas it performed faithful recovery more frequently for large errors than small errors. As in Section 5, there is a significant difference between the effects of error size of faithful and unfaithful recovery rates ($p<0.001$). Similarly, this dissociation between faithful recovery and complete hallucination responses in this experiment also hold for Claude-3 Opus ($p<0.05$) and Llama-3 70B ($p<0.001$).
D.2 Experiment 2
In experiment 2, we measured the effect of context noise on error recovery rates. The trends of the complete hallucinations are largely similar to those of the unfaithful responses presented in the primary analysis. For Claude-3 Opus, we identify a significant difference between faithful recovery and complete hallucination in response to both context noise ($p<0.001$) and the error recovery prompt ($p<0.001$). For Llama-3 70B, we find similar dissociations ($p<0.05$ for context noise and $p<0.001$ for prompting). However, under this sensitivity analysis, we were unable to detect a similarly significant differentiation for GPT-4’s faithful recovery and complete hallucination behaviors ($p>0.05$ for both experimental comparisons).
Figure 18: Sensitivity analysis for faithful vs. unfaithful recoveries, examining the difference in recovery rates between context noise and baseline conditions. Negative values indicate that recoveries occurred more often in the baseline condition. Error bars indicate 95% confidence intervals.
D.3 Experiment 3
@@
Figure 19 shows the complete hallucination rates against the faithful recovery rates for experiment 3. Consistent with our analysis from Section 7, we find that complete hallucination shows distinct behavior compared to faithful recovery (and the overall error recovery trend), from copy errors to calculation errors ($p<0.005$) or from calculation errors to propagated calculation errors ($p<0.001$).